Section 7

Lie Groups and Lie Algebras

In case of several authors, Invited Speakers are marked with a ^{∗}.

James Arthur: Towards a Stable Trace Formula . . . . . . . . . . . . . . . . . . . . II 507

Joseph Bernstein: Analytic Structures on Representation Spaces of Reductive Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 519

Ivan Cherednik: From Double Hecke Algebra to Analysis . . . . . . . . . . II 527
Alex Eskin: Counting Problems and Semisimple Groups . . . . . . . . . . . . II 539

Robert E. Kottwitz: Harmonic Analysis on Semisimple p-Adic Lie Algebras . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 553

L. Lafforgue: Chtoucas de Drinfeld et Applications . . . . . . . . . . . . . . . . II 563
Shahar Mozes: Products of Trees, Lattices and Simple Groups . . . . . II 571

Vera Serganova: Characters of Irreducible Representations of Simple Lie Superalgebras . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 583

Kari Vilonen: Topological Methods in Representation Theory . . . . . . II 595

Minoru Wakimoto: Representation Theory of Affine Superalgebras at the Critical Level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 605

Documenta Mathematica · Extra Volume ICM 1998 · II ·

506

Documenta Mathematica · Extra Volume ICM 1998 · II ·